OpenStudy (anonymous):

Use the normal distribution of SAT critical reading scores for which the mean is 515 and the standard deviation is 118. Assume the variable x is normally distributed. a)What part of the SAT verbal scores is less than 625?

jimthompson5910 (jim_thompson5910):

do you have a TI calculator?

OpenStudy (anonymous):

It really depends on what you are using to calculate the cumulative distribution function.

OpenStudy (anonymous):

If you are using Z-tables \(\Phi(z)\) then \[ z = \frac{X-\mu}{\sigma} \]

OpenStudy (anonymous):

\[ z = \frac{625-515}{118}\\ \Pr(X\leq625) = \Phi\left(\frac{625-515}{118}\right) \]